OpenStudy (anonymous):

Is x-3 a solution to the equation -3x^2=-27?Explain

OpenStudy (lxelle):

-3x^2+27=0

OpenStudy (lxelle):

You will get (x-3)(x+3)=0 So?

OpenStudy (anonymous):

so it is a solutiion?

OpenStudy (anonymous):

@LXelle

OpenStudy (lxelle):

isn't it obvious enough?

OpenStudy (anonymous):

Sorry.Thanks

OpenStudy (anonymous):

May you exlain how you got this awnser as far as solving it???

OpenStudy (anonymous):

Do you know how to solve quadratics ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@AngusV

OpenStudy (lxelle):

From -3x^2 +27 = 0 you deduce it to -x^2 +9 = 0

OpenStudy (anonymous):

And after that, you solve the quadratic equation. Note that in this case a=-1, b=0 and c=9.

OpenStudy (anonymous):

As for solving quadratics, http://lmgtfy.com/?q=how+to+solve+quadratics